OpenStudy (anonymous):

(6-(square root 5)(5+(square root)5

OpenStudy (anonymous):

\[1-5\sqrt{5}\]

OpenStudy (anonymous):

\[(6-\sqrt{5})*(5+\sqrt{5})=30+6\sqrt{5}-5\sqrt{5}-5=25+\sqrt{5}\]

OpenStudy (anonymous):

its a 13 somewhere in the answer

OpenStudy (anonymous):

your right thank you

OpenStudy (anonymous):

the brackets well positioned